please factor by grouping
x-y-x^2+2xy-y2

Answer by stanbon(75887)

please factor by grouping
x-y-x^2+2xy-y2
=(x-y)-(x^2-2xy+y^2
=(x-y)-(x-y)^2
=(x-y)[1-(x-y)]
=(x-y)(1-x+y)
Cheers,
Stan H.